Let u = (4,2,-1), v = (3, 1, 1) and w = (0, 2, 1). Compute the following:
(i) 2v - 3w -u
(ii) u(w+v)
(iii) ||u . w ||
(iv) the orthogonal projection of u on w
(v) the vector component of u orthogonal to w
1
Expert's answer
2020-05-10T15:37:04-0400

"\bold i) \\2v - 3w - u = (6, 2, 2) - (0, 6, 3) - (4, 2, -1) = (2, -6, 0)\\

\bold{ii)} \\
w + v = (3,3,2)
\\u(w + v) = (4,2,-1)(3,3,2) =4*3 +3*2 +2*(-1) = 16 \\

\bold{iv)} \\ proj_w u=\frac{u•w} {|w| }*w\\
Find\ the\ dot\ product\ of u \ and \ w\\
u•w=4•0+2•2+1•(-1)=3\\
Find\ the \ length \ of \ w\\
|w|= \sqrt{0^2+2^2+1^2}=\sqrt{5}\\
proj_wu=(0,\frac{6\sqrt{5}} {5},\frac{3\sqrt
{5}}{5})\\
\bold{v)}\\
comp_w u = u -\frac{u•w}{|w|}w=(4,2,-1) - (0,\frac{6\sqrt{5}}{5},\frac{3\sqrt{5}}{5})=\\
=(4, \frac{10-6\sqrt{5}}{5}, \frac{-5-3\sqrt{5}}{5})\\
\bold{iii)}

\begin{Vmatrix}
 u*w
\end{Vmatrix}
= \begin{Vmatrix}
 i & j & k\\
 4 & 2 & -1\\
 0 & 2 & 1
\end{Vmatrix}
= \begin{Vmatrix}
 \begin{vmatrix}
 2 & -1 \\
 2 & 1\\
\end{vmatrix},&
 \begin{vmatrix}
 -1 & 4 \\
 1 & 0
\end{vmatrix}, 
\begin{vmatrix}
 4 & 2 \\
 0 & 2
\end{vmatrix}\\
 
\end{Vmatrix}
=\begin{Vmatrix}
 (4, -4 , 8)
 
\end{Vmatrix}
= \sqrt{4^2+(-4)^2+8^2} = 4\sqrt{6}"